OpenStudy (anonymous):

Find the unknown value in the percent proportion. part/whole=percent/100. part 30 and whole 300

OpenStudy (anonymous):

@ganeshie8

OpenStudy (ranga):

part/whole = percent / 100 Substitute all the known quantities: part = 30, whole = 300 30 / 300 = percent / 100 Can you now find the percent?

OpenStudy (anonymous):

no i dont know how to.

OpenStudy (ranga):

I can write the above equation as: percent / 100 = 30 / 300 Multiply both sides by 100 and you get: percent = 30 / 300 x 100 = 10%

OpenStudy (anonymous):

thank you

OpenStudy (ranga):

you are welcome.